A traffic engineer records a sample of the types of vehicles that cross a bridge. she counts 112 passenger cars, 18 light trucks, and 20 heavy trucks. she wants to make a spinner to model the probability of each type of vehicle crossing the bridge. Determine the percent of each sector of the spinner to the nearest percent. To the nearest tenth of a percent, what is the probability that at a given moment, a light truck crosses the bridge followed by a heavy truck?


1
Expert's answer
2021-06-03T17:18:34-0400

"112 + 18 + 20 = 150" total vehicles


"\frac{112}{150 }= 0.746 = 74.6\%" rounds to 75% passenger cars

"\frac{18}{150} = 0.12 = 12\%" light trucks

"\frac{20}{150} = 0.133 = 13.3\%" rounds to 13% heavy trucks


P(light truck followed by heavy truck)"= \frac{18}{150} \times \frac{20}{149} =\frac { 12}{745 }" or 1.6%